How many three-digit numbers can be formed under the following conditions?
(a) The leading digit cannot be zero
(b) the leading digit cannot be zero and no repetition of digits
(c) the leading digit cannot be zero and the number must be a multiple of 5
(d) the number is at least 400
for (a) i tried doing 9^3 which gives me 729, the ans. in the back of the book says 900, i also tried doing 9 nCr 3 which gave me 85, and 9!/3!

Answer by checkley77(12844)

(a) The leading digit cannot be zero
The first digit is 1 of 8 the next 2 are 1 of 10.
9*10*10=900 ans.
(b) the leading digit cannot be zero and no repetition of digits
9*9*8=648 ans.
(c) the leading digit cannot be zero and the number must be a multiple of 5
9*8*1=72 ans.
(d) the number is at least 400
1*9*8=72 ans.